### v2

# The role of equation of state on the spin-up of millisecond pulsars

Abstract
Millisecond pulsars (MSPs) are recycled pulsars which have been spun-up due to mass accretion during the phase of mass exchange in binaries. Although the interactions with companion stars play important roles on the spin-up process, the global properties of pulsars determined by the equation of state (EoS), such as mass, radius and the moment of inertia, should also play a role. We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength. The magnetic field strength will be important in impeding the spin-up of NSs, especially to the ones with low masses, but the impeding effect would not be important to SSs. The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS. Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on magnetic field strength, could help to put more strict constraints on the EoS of pulsars.
dense matter — stars: neutron — pulsars: general
1 Introduction
The pulsar-like compact stars are idea laboratories to explore the state of matter at supra-nuclear densities. The equation of state (EoS) of pulsars is still a controversial topic at present, although abundant observational data of pulsars has been accumulated. Based on different points of view, a variety models for pulsars have been speculated. The neutron star (NS) model whose basic idea could be found in Landau (1932) and the quark star (QS) model which originated from Witten’s conjecture (Bodmer, 1971; Witten, 1984) have attracted most attention. The strangeon star (SS) model was also proposed (Xu, 2003) to understand different manifestations of pulsars, where the strong coupling between quarks has grouped quarks into strangeons (i.e. “strange nucleons”) , and has passed various observational tests (see Lai et al. (2023) and references therein).
The EoS of pulsars is essentially depends on the behavior of quantum chromodynamics (QCD) at low-energy scales, which is still a challenge for us to understand. Although the constraints by both the existence of high-mass pulsars and the tidal deformability of GW 170817 (Abbott et al., 2017) could dramatically reduce the allowed EoSs of NSs (Annala et al., 2018), it is still uncertain that whether pulsars are NSs, QSs or SSs. In addition, these three kinds of models mentioned above have subclass models respectively, e.g., some branches of NS model still allow the existence of quark matter at the core region (Weber, 2005), QS model includes different states such us unpaired strange quark matter state (Alcock et al., 1986) and colour super-conductivity (CSC) state (Alford et al., 2008), and the hybrid SS model is proposed (Zhang et al., 2023) where an SS develops a strange quark matter core. Therefore, the differences among models could be subtle. It is certainly important to find out unambiguous and observable probes to tell what is the real nature of pulsars.
The global structure of pulsars, e.g. the mass and radius, can provide robust constraints on EoS models when compared with observations (Lattimer & Prakash, 2007; Özel et al., 2010). Despite the uncertainty about the explicit structure determined by various classes of EoS models, there are some general differences in the global structures between NSs and QSs/SSs (Demorest et al., 2010; Gao et al., 2022). NSs are gravity-bound so that the surface density could approach zero, and the radii of low-mass NSs (below $\sim M_{\odot}$) decrease with increasing masses and the radii of intermediate-mass NSs (in the range between $\sim M_{\odot}$ and $\sim 2M_{\odot}$) could remain nearly constant. Both QSs and SSs are self-bound by strong interaction so that the surface density could still above the nuclear matter density, and their radii increase with increasing masses unless the mass approaches its maximum value. The distinctions due to gravity-binding and self-binding, especially for the mass below $\sim 1.5M_{\odot}$ when QSs/SSs would be more compact than NSs with the same mass, seem to be more pronounced than that due to, e.g., whether the degrees of freedom are free quarks or strangeons, or whether quark matter emerges inside NSs or not. Therefore, testing whether pulsars are gravity-bound or self-bound provides a relatively less unambiguous probe to the nature of pulsars.
In this paper we investigate that how different global structures of pulsars from different EoS models affect the spin-up process of millisecond pulsars (MSPs). MSPs evolve from pulsars in binaries which have underwent the spin-up process, i.e. recycling precess, due to accreting mass from their companions. The formation of MSPs with a carbon-oxygen white dwarf (CO WD) (Tauris et al., 2000, 2011) or a helium white dwarf (He WD) (Tauris et al., 2012; Chen et al., 2021) in intermediate-mass X-ray binaries (IMXBs) and low-mass X-ray binaries (LMXBs) (Podsiadlowski et al., 2002) have been investigated to show that the spin-period of a recycled MSP is related to the amount of accreted mass which depends on the nature of the donor star. The observed differences of the spin distributions of recycled pulsars with different types of companions have been explained by the amount of accreted mass needed for a pulsar to achieve its equilibrium spin (Tauris et al., 2012). In addition, the magnetic field evolution of accreting pulsars would also affect the formation of MSPs. The spin-period evolutions of recycled pulsars could have been influenced by the accretion induced magnetic field decay (Cheng & Zhang, 1998; Zhang & Kojima, 2006). The minimum spin-period of millisecond pulsars also depends the correlation between the accretion rate and the surface dipole magnetic field in LMXBs (Ertan & Alpar, 2021).
The spin-up line in the $P\dot{P}$-diagram is often used to demonstrate the formation and evolution of recycled pulsars, although it in fact cannot be uniquely defined (Tauris et al., 2012) due to the dependence on both the process of accretion and the properties of the pulsar. Most of the previous studies about the evolution of MSPs focusing on the accretion process in the binary evolution use the typical values for the structure of neutron stars. These studies discuss many parameters which do affect the location of the spin-up line in the $P\dot{P}$-diagram, e.g. the critical fastness parameter (Wang et al., 2011) and the location of magnetospheric boundary (Ghosh & Lamb, 1979a). It is worth noting that, the spin evolution of a pulsar also depends on the EoS which determines the pulsar’s global structure such as mass, radius and the moment of inertia. In this paper we will investigate the role of equation of state on the spin-up of MSPs.
Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs. From the spin-up condition we derive the spin-up lines in $MP$ (mass vs spin-period) diagram in Sec.2, and find that under the same magnetic field strength an SS could spin faster than an NS in the recycling process. To take into account the effect of accretion process, in Sec.3 we derive the minimal spin-period by assuming that all the angular momentum of the accreted material is added the the pulsar. Although the results of Sec.3 show that NSs would spin faster than SSs with the same amount of accreted mass and the same magnetic field, the minimal-spin contours derived in Sec.4 show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same. The results shown in Sec.4 indicate that the constraints on EoS models are significant for masses below about $1.5M_{\odot}$, especially when we get more accurate estimation of the magnetic field strength. Conclusions and discussions are made in Sec.5.
2 The spin-up line
[…]
Due to the model-dependence of the accretion torque on the spin-period and the magnetic field configuration (Ghosh & Lamb, 1979a), to be spun-up by the accretion the pulsar should have the spin angular velocity $\Omega$ satisfying $\Omega\leq\omega_{\rm c}\cdot\Omega_{\rm K}(r_{\rm m})$, where $\Omega_{\rm K}(r_{\rm m})$ is the Keplerian angular velocity at radius $r_{\rm m}$, and $\omega_{\rm c}$ is the critical fastness parameter. To exert a positive accretion torque, $\omega_{\rm c}$ is about 0.2-1 (Ghosh & Lamb, 1979b; Wang et al., 2011).
The spin-up condition can therefore be expressed as the relation between $r_{\rm A}$ and $r_{\rm co}$, where the corotation radius $r_{\rm co}=[GMP^{2}/(4\pi^{2})]^{1/3}$ is defined as the radius where the Keplerian frequency matches the spin frequency of the pulsar. Using the factors $\phi$ and $\omega_{\rm c}$ mentioned before, the spin-up condition can be written as $r_{\rm A}\leq[\omega_{\rm c}^{2/3}/\phi]r_{\rm co}$. Because we will focus on the EoS-dependent results of the spin-up process, the dependence on other factors, such as truncation of the accretion disc and the magnetic field configuration outside the star, will not be included when we are comparing different EoS models. Taking into account the possible values for the factors $\phi$ and $\omega_{\rm c}$ mentioned before, we assume $\omega_{\rm c}^{2/3}/\phi=0.5$, so the spin-up condition is
$r_{\rm A}\leq 0.5\ r_{\rm co}.$ (1)
The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar). However, the real spin-down torque should be more complicated than the magnetic dipole braking in vacuum. For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006). To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities. Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram.
The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent. Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.1 shows the results for NS and SS models. Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively. We can see that the spin-up lines of NSs are decreasing with the mass, and that of SSs are nearly independent of the mass. The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
Figure 1: Spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
3 The minimal spin-period by accretion
By accreting the amount of mass ${\rm d}M$, the spin angular momentum added to a pulsar is given by ${\rm d}J=n\xi\sqrt{GMr_{\rm A}}{\rm d}M$, where $\xi\simeq 1$ is a numerical factor depending on the flow pattern (Ghosh & Lamb, 1979a) and $n$ is a dimensionless torque which could be assumed to be $\simeq 1$ (Tauris et al., 2012). Then the final spin-period $P$ of a recycled pulsar (from rest) from the initial mass $M_{1}$ to the final mass $M_{2}$ could be estimated by
$P^{-1}_{\rm min}=\frac{1}{2\pi I}\int_{M_{1}}^{M_{2}}\sqrt{GMr_{\rm A}}{\rm d}M,$ (2)
where $I$ is the (final) moment of inertia. We have assumed above that all the orbital angular momentum of the accreted material is transferred to the spin angular momentum of the pulsar, so $P$ in equation (2) is the minimal value of the final spin-period. In the following the values of $I$ are derived when the spin angular momentum of the star is calculated to the first order of spin angular velocity $\Omega$ (Gao et al., 2022), which means that $I$ is independent of $\Omega$, and this approximation is good when the $P>1$ ms.
Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables. To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumption as in Tauris et al. (2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium (but this assumption can be relaxed in deriving the minimal-spin contour which contains different values of $B$, see Sec.4). In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak. Consequently, the variables in the right-hand side of equation (2) are $M_{1}$, $M_{2}$, $M$, $R$ and $I$. Given an EoS, the minimal spin-period of a recycled pulsar with the amount of accreted mass $\Delta M=M_{2}-M_{1}$ could be derived.
The $P_{\rm min}$ and $M$ (final mass) relation for SS and NS models are shown in Fig.2, denoted respectively by red lines and blue lines, where solid lines represent the case of $\Delta M=0.05M_{\odot}$ and dashed lines represent the case of $\Delta M=0.1M_{\odot}$. We can see that, with the same amount of accreted mass, NSs would spin faster than SSs. This is understandable: the radius of an NS is usually larger than that of an SS with the same mass, so an NS usually has larger $r_{\rm A}$ than an SSs, which means that an NS could gain more angular momentum than an SS when the accreted mass is the same. In Fig.2 we choose $B=5\times 10^{8}$ G, and smaller values of $B$ will lift the lines upwards, since smaller $r_{\rm A}$ will lead to smaller angular momentum transfer.
Figure 2: The minimal spin-period by accreting the amount of mass $\Delta M$ for both SS (red lines) and NS models (blue lines), when the strength of magnetic field $B=5\times 10^{8}$ G. The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively. In $MP$ diagram, the horizontal ordinate denotes the final mass after accretion.
4 Comparison with observations
The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram. As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$. Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$.
If the values of both $B$ and $\Delta M$ of a recycled pulsar is known, we can plot the allowed region in $MP$ diagram which satisfies both equations (1) and (2), i.e. by putting Fig.1 and Fig.2 together. We show the allowed regions of SSs (left) and NSs (right) for $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G in Fig.3, where dashed line denotes the general spin-up line (cf. Fig.1) and solid line shows the minimal spin-period by accretion (cf. Fig.2). From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
Refer to captionRefer to caption
Figure 3: The allowed regions of SSs (left) and NSs (right) in $MP$ diagram, where solid curves denote minimal-spin lines with $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G, and dashed curves denote spin-up lines with $B=5\times 10^{8}$ G.
However, in most cases we are not certain about both the value of $B$ and $\Delta M$ for recycled pulsars. If we estimate $B$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, then by setting $\sin\alpha=1$ we can get $B>1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$. It is interesting to noted that, each pulsar whose data of $M$ and $P$ is plotted in Fig.5 (see below) is above the spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue (https://www.atnf.csiro.au/research/pulsar/psrcat/(Manchester et al., 2005) under both SS and NS models. As mentioned before, in this paper we keep $B$ as a parameter, instead of replacing in some model-dependent ways. Then how to show the difference between SSs and NSs models by comparing with observations?
Combining equations (1) and (2) we can get the minimal-spin contour by accretion of $\Delta M$, which is defined as the contour of $\Delta M$ in $MP$ diagram, representing the minimal spin-period by accretion of $\Delta M$ for different values of $B$. For a data point in $MP$ diagram, there exists both a minimal-spin contour and a spin-up line intersecting it. The value of $\Delta M$ of the minimal-spin contour is the minimal accreted mass of the pulsar, while the value of $B$ of the spin-up line is the upper limit on its magnetic field strength. If we can use some other methods to constrain $\Delta M$ (e.g. from the binary evolution) and $B$ (e.g. theoretically from the magnetic field decay in accretion, or observationally from the cyclotron absorption lines), it would be hopeful to see that whether this pulsar is more likely to be an NS or an SS.
From Fig.3 we can infer that low mass and fast spinning pulsars would present significant different parameters between SS and NS models, so we use a hypothetical data point in $MP$ diagram as an example in Fig.4. If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G. One the other hand, if this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G. This example shows that the difference in the the upper limit on $B$ between SS and NS models is significant for low mass and fast spinning pulsars.
Figure 4: We use a hypothetical data point in $MP$ diagram as an example. If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G. If this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
To compare with actual data, we extend the curves in Fig.4 to different values of $B$ and $\Delta M$, and we show in Fig.5 the results for SSs (left) and NSs (right). Solid curves correspond to minimal-spin contours with $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to spin-up lines with $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
Refer to captionRefer to caption
Figure 5: Minimal-spin contours of $\Delta M$ and spin-up lines of $B$ for SSs (left) and NSs (right). The horizontal ordinate denotes the final mass after accretion. Solid curves correspond to $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B/10^{8}{\rm G}=2,4,8,12$, increasing upwards. Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.5. Although the minimal-spin contours of $\Delta M$ for SSs are not significantly different from that for NSs, the differences between the contours of $B$, i.e. the spin-up lines, for SSs and that for NSs are visible. The dashed curves in Fig. 5 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs. The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$ than that for higher masses.
If we can find some fast spinning pulsars with low masses, and if we can additionally know the lower limit of $B$, i.e. $B>B_{0}$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS. For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could. As we mentioned before, each pulsar plotted in Fig.5 is above its spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using its data of $P$ and $\dot{P}$, so at this point if we take the lower limit magnetic field strength $B_{0}$ as $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ we could not yet disfavor neither SS nor NS models. Certainly, finding more MSPs of $P<10$ ms with low mass and more accurate estimation of $B_{0}$ in the future could help to put more strict constraints on the EoS of pulsars.
5 Conclusions and discussions
The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries. We investigate the spin-up of MSPs in AP4 EoS of NS model and LX3630 EoS of the SS model. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs. Under the same magnetic field strength, an SS could spin faster than an NS in the recycling process, and NSs would spin faster than SSs with the same amount of accreted mass. The minimal-spin contours show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same, but the differences between the spin-up lines for SSs and that for NSs are obvious. The magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs, so it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$, when the upper limit of the allowed magnetic field strength of an SS is higher than that of an NS. To give a more general and model-independent investigation, in this paper we keep the magnetic field strength $B$ as a parameter instead of replacing it in some model-dependent ways. It is worth noting that an accreting pulsar would have an lower limit of magnetic field strength (e.g. see Zhang & Kojima (2006)), so the constraints on NS model would be more strict than that on SS model. For the fast spinning pulsars with low masses, if we can additionally derive the lower limit of magnetic field strength theoretically from the magnetic field decay models, or if we can directly detect the magnetic field strength from the cyclotron absorption lines, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km. We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs. To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as the pattern of disk-fed accretion flow and the magnetic field configuration outside the accreting star, which lead some simplifications in equations (1) and (2). In addition we do not consider the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion. Still, our results indicate that it is worth adding the role of EoS to other effects to study the recycling of MSPs more comprehensively.
The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions. A gravity-bound pulsar, as shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, as shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value. This difference leads to different evolutions of $r_{\rm A}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars. We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star. Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on magnetic field strength in the future, e.g. by SKA and FAST, could be helpful to put more strict constraints on the nature of pulsars.
This work is supported by the National SKA Program of China (No. 2020SKA0120300) and the Young Top-notch Talent Cultivation Program of Hubei Province.